# Cite Before You Speak: Enhancing Context-Response Grounding in E-commerce Conversational LLM-Agents

### v2

@@ Abstract
With the advancement of conversational large language models (LLMs), several LLM-based Conversational Shopping Agents (CSA) have been developed to help customers answer questions and smooth their shopping journey in e-commerce domain. The primary objective in building a trustworthy CSA is to ensure the agent’s responses are accurate and factually grounded, which is essential for building customer trust and encouraging continuous engagement. However, two challenges remain. First, LLMs produce hallucinated or unsupported claims. Such inaccuracies risk spreading misinformation and diminishing customer trust. Second, without providing knowledge source attribution in CSA response, customers struggle to verify LLM-generated information. To address these challenges, we present an easily productionized solution that enables a “citation experience” utilizing In-context Learning (ICL) and Multi-UX-Inference (MUI) to generate responses with citations to attribute its original sources without interfering other existing UX features. With proper UX design, these citation marks can be linked to the related product information and display the source to our customers. In this work, we also build auto-metrics and scalable benchmarks to holistically evaluate LLM’s grounding and attribution capabilities. Our experiments demonstrate that incorporating this citation generation paradigm can substantially enhance the grounding of LLM responses by 13.83% on the real-world data. As such, our solution not only addresses the immediate challenges of LLM grounding issues but also adds transparency to conversational AI.
@@ 1 Introduction
Figure 1: An Overview of Citation Generation Paradigm. For this specific user query, Customer Reviews are used as in-context evidences to generate a more accurate and reliable response with easily verifiable knowledge sources.
@@
However, accurately answering customer questions with precise product information can be challenging. Firstly, for a given customer query and product, finding correct and reliable product facts is non-trivial. Product descriptions provided by sellers are sometimes inconsistent with actual customer experiences. For example, shoes advertised as “true to size” might turn out to be “oversized” for the majority of customers. Secondly, even when reliable information is available, LLMs may still hallucinate [1, 2]. Retrieval-augmented generation (RAG) [3, 4, 5] can alleviate the hallucination problem in LLMs by augmenting retrieved information directly into the generation process, however, its effectiveness highly depends on the relevancy of the retrieval sources [6]. Our analysis reveals that approximately $3\%$ of claims in CSA-generated responses are not grounded on the product facts provided. When customers verify the sources of the information and discover discrepancies themselves, it will significantly reduce customer-CSA trust, resulting in diminished retention on CSAs over the long run.
In this light, we propose a novel framework to enable “citation experience” on CSA that links LLM-generated responses to reliable product facts. In-context Learning (ICL) has been demonstrated as the emerging capability of LLMs [7, 8, 9]. Research has shown the effectiveness of using ICL to instruct LLMs to generate responses while providing citations of the origins in the form of box brackets (i.e. $[n]$) [10]. However, one of the biggest challenges in productionizing the citation generation paradigm as a UX feature on CSAs via ICL is that without explicit model training to enhance the backbone LLM’s attribution capability, the quality of the generated citations might be compromised and may interfere with other UX features that have been trained and deployed. To ensure both quality and efficient latency to build a Minimum Viable Product (MVP), we propose a solution utilizing ICL accompanied by a Multi-UX Inference (MUI) technique to address the aforementioned issues at decoding time. Our approach comprises three key components:
  • Easily Productionized Solution for Knowledge Source Attribution in LLM-Generated Responses: Inspired by LLM-based search systems like Perplexity or Gemini, we implement an easily productionizing solution via ICL and paged attention [11] to enable LLM citing information sources directly in LLM responses. By appending citations to statements, customers can easily trace information back to verified product reviews and descriptions.
  • Leveraging Real Customer Reviews: We feed customer reviews into the generative LLM to mitigate inaccuracies stemming from unreliable seller-provided marketing descriptions. The reviews reflect true customer feedback on the products, offering a more reliable source of product information.
@@
In this work, our key contributions are: ➊ Propose an easy-built solution to enable citation generation on CSA, which is designed to be easily moved into production as a Minimum Viable Product (MVP) to gather customer feedback ➋ Propose a MUI to address the productionization challenge of different UX features potentially interfering with each other’s performance ➌ Discover insights about citation generation paradigm on improving response grounding.
@@ 2.1 Problem Formulation
In retrieval-augmented generation, for a given user query $q$, the retriever first retrieves a set of evidences $\mathcal{E}$. A generative LLM, $\mathcal{M}$ takes the prompt, which consists a user query $q$, the retrieved evidence set $\mathcal{E}$, and a system prompt $P$ to produce a response $\mathcal{R}=\mathcal{M}(q,\mathbb{E},p)$. The response =$\mathcal{R}=\{s_{1},s_{2},...,s_{n}\}$ consists of $n$ sentences. Each sentence $s$ cites from a list of evidences $\mathcal{E}^{s}=\{e_{1}^{s},...,e_{l}^{s}\}\subseteq\mathcal{E}$. If a sentence is generated without any citations, $\mathcal{E}^{s}=\emptyset$.
@@ 2.2 Citation Generation Methodology
With the increasing capabilities of large language models (LLMs), in-context learning (ICL) has been demonstrated as an emerging ability of LLMs to follow instructions and learn from a few examples in the in-context [9]. In-context learning is a powerful and convenient approach for adapting LLMs for downstream tasks. Compared with other approaches, ICL provides an efficient way to build Minimum Viable Products (MVPs) and test user feedback. Post-hoc approach [12] adopts Natural Language Inference (NLI) models to capture the entailment relations between each evidence and each sentence in CSA’s response, and add the evidence citations to the response if the evidence supports the sentence. However, adding citations retroactively can significantly increases the latency of the response generation. Fine-tuning approach [12, 13, 14], on the other hand, can further adapt the pre-trained LLMs to generate responses with attributable sources. However, it requires a holistic training approach to integrate attributable information generation into the answer generation process without degrading the model’s other fundamental capabilities. Additionally, model training requires high-quality data to be collected and annotated, which is time-consuming and labor-intensive. With all these considerations, we choose the ICL-based approach [10] to enable citation generation.
@@ 3.3 Evaluating LLM’s Evidence Utilization
4 Experiment Setup and Results
@@
5 Multi-UX Inference (MUI)
The CSA production system not only answers product factoid questions but also supports other user experiences (UXs), such as related question suggestions. To avoid multiple LLM inference calls for the same customer request, one single prompt is used to generate the response for all UXs. To enable citation functionality in CSA responses, the corresponding instructions need to be appended to this prompt. However, our analysis shows that adding citation instructions to the prompt leads to performance regressions in other UXs. To address this issue without using multi-inference calls, we propose leveraging paged attention [11] to cache the shared prompt. Specifically, when introducing a new UX via ICL, the corresponding instructions are appended to the end of the original prompt. During the prefilling stage, we compute the key-value (KV) cache for the entire prompt. In the decoding stage, we use the full prompt to generate responses with citations, while excluding the citation instructions when generating responses for other UXs. The KV cache corresponding to the citation instructions is stored on a separate page. To prevent citation instructions from negatively affecting existing UXs, we simply avoid fetching that page during decoding. This approach ensures that we maintain high performance across all UXs while enabling citation functionality in the same prompt.

### v3

@@ Abstract
With the advancement of conversational large language models (LLMs), several LLM-based Conversational Shopping Agents (CSA) have been developed to help customers smooth their online shopping. The primary objective in building an engaging and trustworthy CSA is to ensure the agent’s responses about product factoids are accurate and factually grounded. However, two challenges remain. First, LLMs produce hallucinated or unsupported claims. Such inaccuracies risk spreading misinformation and diminishing customer trust. Second, without providing knowledge source attribution in CSA response, customers struggle to verify LLM-generated information. To address both challenges, we present an easily productionized solution that enables a “citation experience” to our customers. We build auto-evaluation metrics to holistically evaluate LLM’s grounding and attribution capabilities, suggesting that citation generation paradigm substantially improves grounding performance by 13.83%. To deploy this capability at scale, we introduce Multi-UX-Inference system, which appends source citations to LLM outputs while preserving existing user experience features and supporting scalable inference. Large-scale online A/B tests show that grounded CSA responses improves customer engagement by 3%10%, depending on UX variations.
**footnotetext: Equal Contribution.
@@ 1 Introduction
Figure 1: An Overview of Citation Generation Paradigm. For this specific online customer query, Customer Reviews are retrieved as in-context evidences to generate a more accurate and reliable response with easily verifiable knowledge sources.
@@
However, accurately answering customer questions with precise product information can be challenging. Firstly, for a given customer query and product, finding correct and reliable product facts is non-trivial. Product descriptions provided by sellers are sometimes inconsistent with actual customer experiences. For example, shoes advertised as “true to size” might turn out to be “oversized” for the majority of customers. Secondly, even when reliable information is available, LLMs may still hallucinate [1, 2]. Retrieval-augmented generation (RAG) [3, 4, 5, 6] can alleviate the hallucination problem in LLMs by augmenting retrieved information directly into the generation process, however, its effectiveness highly depends on the relevancy of the retrieval sources [7]. Our analysis reveals that approximately $3\%$ of claims in CSA-generated responses are not grounded on the product facts provided. When customers verify the sources of the information and discover discrepancies themselves, it will significantly reduce customer-CSA trust, resulting in diminished retention on CSAs over the long run.
In this light, we propose a novel framework to enable “citation experience” on CSA that links LLM-generated responses to reliable product facts. In-context Learning (ICL) has been demonstrated as the emerging capability of LLMs [8, 9, 10, 11]. Research has shown the effectiveness of using ICL to instruct LLMs to generate responses while providing citations of the origins in the form of box brackets (i.e. $[n]$) [12]. However, one of the biggest challenges in productionizing the citation generation paradigm as a UX feature on CSAs via ICL is that without explicit model training to enhance the backbone LLM’s attribution capability, the quality of the generated citations might be compromised and may interfere with other UX features that have been trained and deployed. To ensure both quality and efficient latency to build a Minimum Viable Product (MVP), we propose a solution utilizing ICL accompanied by Multi-UX Inference (MUI) technique to address the aforementioned issues at decoding time. Our approach comprises three key components:
  • Easily Productionized Solution for Knowledge Source Attribution in LLM-Generated Responses: Inspired by LLM-based search systems like Perplexity or Gemini, we implement an easily productionizing solution via ICL to enable LLM citing information sources directly in LLM responses. By appending citations to statements, customers can easily trace information back to verified product reviews and descriptions. We also address inference challenges through MUI with inference optimization techniques.
  • Leveraging multi-perspective product information: we dynamically supply the generative LLMs with multi-perspective in-context content, such as merchants’ product descriptions, customer reviews, customer Q&A, and etc., to let the CSA generate more well-rounded, reliable answers. The multi-perspective knowledge source provided as in-context can mitigate inaccuracies stemming from unreliable seller-provided marketing descriptions.
@@
In this work, our key contributions are: ➊ Propose a lightweight citation-generation solution, paired with an automated, holistic evaluation framework—both engineered for rapid MVP deployment to gather real-world customer feedback. ➋ Introduce Multi-UX Inference (MUI), a technique that prevents ICL-based UX features from degrading one another’s performance; and an efficient request orchestration layer that scales LLM inference ➌ Demonstrated impacts of citation generation paradigm on improving response grounding by 13.83% and customer engagement rate by 3-10%.
@@ 2.1 Problem Formulation
In retrieval-augmented generation, for a given online customer query $q$, the retriever first retrieves a set of evidences $\mathcal{E}$. A generative LLM, $\mathcal{M}$ takes the prompt, which consists an online customer query $q$, the retrieved evidence set $\mathcal{E}$, and a system prompt $P$ to produce a response $\mathcal{R}=\mathcal{M}(q,\mathbb{E},p)$. The response =$\mathcal{R}=\{s_{1},s_{2},...,s_{n}\}$ consists of $n$ sentences. Each sentence $s$ cites from a list of evidences $\mathcal{E}^{s}=\{e_{1}^{s},...,e_{l}^{s}\}\subseteq\mathcal{E}$. If a sentence is generated without any citations, $\mathcal{E}^{s}=\emptyset$.
@@ 2.2 Citation Generation Methodology
With the increasing capabilities of large language models (LLMs), in-context learning (ICL) has been demonstrated as an emerging ability of LLMs to follow instructions and learn from a few examples in the in-context [10]. In-context learning is a powerful and convenient approach for adapting LLMs for downstream tasks. Compared with other approaches, ICL provides an efficient way to build Minimum Viable Products (MVPs) and test user feedback. Post-hoc approach [13] adopts Natural Language Inference (NLI) models to capture the entailment relations between each evidence and each sentence in CSA’s response, and add the evidence citations to the response if the evidence supports the sentence. However, adding citations retroactively can significantly increases the latency of the response generation. Fine-tuning approach [13, 14, 15], on the other hand, can further adapt the pre-trained LLMs to generate responses with attributable sources. However, it requires a holistic training approach to integrate attributable information generation into the answer generation process without degrading the model’s other fundamental capabilities. Additionally, model training requires high-quality data to be collected and annotated, which is time-consuming and labor-intensive. With all these considerations, we employ ICL-based approach [12] to enable citation generation and dynamically provide LLM multi-perspective sources—product descriptions, customer reviews, and Q&A as in-context to increase response credibility.
@@ 3.3 Evaluating LLM’s Evidence Utilization
4 Offline Experiment Setups and Evaluation Results
@@
5 Online Deployment and Evaluation
Deployment considerations: The CSA system is deployed in multiple UXs while some UXs such as recommending products do not need citations. The existing product generates the responses to all the UXs using a unified prompt to guarantee the consistency between different UXs. To enable citation experience in production, the most straightforward approach is directly add citation instructions to the unified prompt. However, our analysis shows that this approach leads to performance regressions in other UXs, and LLM could hallucinate to add citations to the UXs that do not need citations. To maintain the answer quality and avoid multiple LLM inference calls for the same customer request, we propose Multi-UX-Inference that leverages paged attention [18] to cache the shared prompt. Specifically, when introducing a new UX via ICL, the corresponding instructions are appended to the end of the original prompt such that the KV cache of the original prompt can be reused. During the prefilling stage, we compute the key-value (KV) cache for the entire prompt. The KV cache corresponding to the citation instructions is stored on separate pages. In the decoding stage, we use the full prompt to generate responses with citations, while excluding the citation instructions page when generating responses for other UXs. We found this approach could maintain comparable answer quality without multiple prefilling, while boosts the inference speed and saves the GPU memory cost.
Figure 2: Multi-UX-inference design with paged-attention.
Online A/B test results: We ran a 2-week online A/B test to evaluate the impact of grounded CSA answers, where the customers are evenly partitioned into control and treatment groups. The large-scale A/B test involves over 10 million customers and shows that improving the system’s answer grounding capabilities can increase customer engagement rate of CSA by 3–10% on different UXs.